# Groups which are metrically weakly sofic with respect to word norms

### v1

@@ 1 Introduction
Let $G$ be a group and let $\Lambda$ be a closed convex subset of $[0,\infty)$ so that $0\in\Lambda$.
@@
If $\ell$ is a $\Lambda$-(pseudo)-norm on $G$ then we say that $(G,\ell)$ is a (pseudo) normed group. The pseudo-norm $\ell$ is invariant if $\ell(h^{-1}gh)=\ell(g)$ for all $g,h\in G$. In this case it defines an invariant pseudo-metric by $d_{\ell}(g,h)=\ell(gh^{-1})$. Thus
@@
There are metric/normed groups for which the question of metric weak soficity looks very interesting. For example it would be interesting to know how it fits to bounded normal generation (with norms as in [7]), in particular in the case of Chevalley groups, see [19]. We also mention locally compact groups with property PL, see [5]. Our original motivation is due to [6], [10], [15].
@@ 1.1 Word norms
Finitely generated groups with word norms are especially visible in (geometric) group theory, see [1], [3], [12], [13], [18] and [21]. For example when $G$ is generated by finitely many conjugacy classes then finiteness of its diameter for the appropriate word metric implies boundedness of its diameter with respect to any invariant metric, [1]. Thus metric weak soficity in the case of word norms is natural for investigations. This leads to an interesting subclass of abstract finitely generated weakly sofic groups: those ones which are metrically weakly sofic with respect to a word norm associated with some tuple of generators. It is an easy exercise that finitely generated abelian groups are metrically weakly sofic with respect to word norms arising from (Maltsev) bases. In fact they are metrically RF (See 2.3). There are other examples in [11] (which are metrically LEF). These results support the following problem:
@@
In order to approach this problem we give a characterisation of metric weak soficity in terms of profinite topology. The following statement is the main result of the paper.
Theorem 4.1. Let $\mathsf{F}$ be a finitely generated free group and $G=\mathsf{F}/N$. Assume that $S\subseteq\mathsf{F}$ is a finite symmetric set representing elements of $G$ which define an invariant word norm, say $\parallel\cdot\parallel_{N}$, in $G$.
The normed group $(G,\parallel\cdot\parallel_{N})$ is metrically weakly sofic if and only if for any finite sequence $g_{1}$, $g_{2}$,…, $g_{k}\in G$ of $\parallel\cdot\parallel_{N}$-norm 1 and $h_{1}$, $h_{2}$,…, $h_{l}$ from $N$ (possibly with repetitions) the closure of the product $g_{i_{1}}^{\mathsf{F}}\cdot g_{i_{2}}^{\mathsf{F}}\ldots\cdot g_{i_{m}}^{\mathsf{F}}\cdot h_{1}^{\mathsf{F}}\cdot h_{2}^{\mathsf{F}}\ldots h_{l}^{\mathsf{F}}$ is a subset of $\{w\in\mathsf{F}|\parallel wN\parallel_{N}\leq m\}N$.
@@
In Sections 3 and 5 we give theorems which characterize metric residual finiteness and metric LEF in terms of profinite topology. We also discuss a number of questions which arise in our approach.
@@ 2 Length functions on a group
When we consider a normed group $(G,\ell)$ it is usually assumed that the norm is invariant. A free group with the standard length function with respect to a free basis will be the only exception of this rule. Just in case in Section 2.1 we do not make the general invariantness assumption.
@@ 2.1 Standard properties of norms
Then the family $\{B_{<r}(g)\,|\,r\in(\Lambda\cap\mathbb{Q})\setminus\{0\},\,g\in G\}$ is a subbase of the topology defined by $\ell$. Thus every normed group is also a topological group.
@@ 2.3 Metric RF and metric LEF
It is worth noting that when in Definition 2.7 the values of $\ell_{G}$ and all possible $\ell_{C}$ belong to $\mathbb{Q}$ (in particular, when $\ell_{C}$ are integer-valued) the definition can be essentially simplified. The condition should be formulated as follows: for every finite $D\subset G$ there is a normed $(C,\ell_{C})\in\mathcal{C}$ and an injective map $\phi:D\rightarrow C$ such that
@@ 2.4 Norms for RF and metric LEF
It is easy to see that $\ell$ is an invariant norm which defines the profinite topology. In order to verify that $(G,\ell)$ is metrically fully residually finite take any finite $K\subset G$. Let $N_{s}$ be chosen so that $K\cap N_{s}\subseteq\{1\}$. By Lemma 2.2 we have $\ell(g)=\ell_{G/N_{s}}(gN_{s})$ for any $g\in K$. ∎
@@
Let $\ell$ be an invariant norm on a group $G$ and $\Lambda_{0}=\mathsf{Rng}(\ell)$. Assume that $(G,\ell)$ is metrically LEF (RF as a metric group).
Then $(G,\ell)$ is metrically LEF (resp. RF as a metric group) with respect to the subclass consisting of finite invariantly normed groups with norms having the set of values in $\Lambda_{0}$.
@@
Take a finite $K\subset G$, a finite $Q\subset\Lambda\cap\mathbb{Q}$ with $0\in Q$ and a $K$-$Q$-almost-homomorphism $\varphi$ from $(G,\ell)$ to some finite $(G_{0},\ell_{0})$ with an invariant norm. It can happen that $\mathsf{Rng}(\ell_{0})\not\subseteq\Lambda_{0}$. In order to correct this let us define a function $\ell^{\prime}$ on $G_{0}$ as follows. If $g\in G_{0}$ and $\ell_{0}(g)\in\Lambda_{0}$ put $\ell^{\prime}(g)=\ell_{0}(g)$.
@@
In order to see property (iii) of Definition 1.1 let us apply induction. Let $m$ be the first number from $\Lambda_{0}$ such that $\ell_{0}$ takes values from $[0,m]\setminus\mathbb{N}$. Thus these reals belong to the open interval $(m-1,m)$. Since all values of $\ell_{0}$ which do not accede $m-1$, belong to $\mathbb{N}$, one easily sees that the (triangle) inequality (iii) holds for $\ell^{\prime}$ and for all $g$, $h$ with $\ell_{0}(gh)\leq m$.
@@
It seems to us that the method used in this proof can work in many other situations. In fact we do not know any example where the natural version of Proposition 2.10 does not hold. For example, it is easy to see that the statement of this proposition holds in the case when $\Lambda_{0}$ is dense in $\Lambda$.
3 Approximations of finitely generated metric groups with respect to word norms
@@ 3.1 Plan for Sections 3,4,5
Our aim is to describe the properties from the problem above in terms of the profinite topology of $\mathsf{F}$. In Section 3 we will study metric RF, in Section 4 - metric weak soficity and in Section 5 - metric LEF. In Section 4 we wiil show that $\mathsf{F}_{2}$ is not metrically weakly sofic as a metric group of the following form.
@@
In the remaining part of this introduction we give two useful remarks.
@@
This remark also holds for residual finiteness of metric groups where $\phi$ to be an abstract homomorphism from $G$ to a finite group.
In the second remark we again arrive at the the situation when $G=\langle\overline{S}\rangle$ with the corresponding word metric but now without assumption of finiteness of $S$. Note that the equalities $B_{1}(1)=\overline{S}\cup\{1\}$ and $B_{n}(1)=(B_{1}(1))^{n}$ always hold in such a group. We now notice that the converse also holds.
@@ 3.2 Metric RF
    The metric group $(G,\parallel\cdot\parallel_{N})$ is RF as a metric group;
@@
    The metric group $(G,\parallel\cdot\parallel_{N})$ is metrically RF;
@@
    For any $m\in\mathbb{N}$ the set $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ is closed in the profinite topology of $\mathsf{F}$.
Then $2\Rightarrow 3\Rightarrow 4\Rightarrow 1$, and assuming that $S$ is finite $1\Leftrightarrow 2$.
@@
$2\Rightarrow 3$. W.l.o.g. we may assume that there are elements $g\in G$ with $\parallel g\parallel_{N}\geq 2$. In order to prove that the set $B^{G}_{m}=\{x\in G\,|\,\parallel x\parallel_{N}\leq m\}$ is closed, take any $g\not\in B^{G}_{m}$. Choose a metric homomorphism $\varphi$ from $G$ to some finite $(H,\ell)$ such that $\ell(\varphi(g))=\parallel g\parallel_{N}$. If $\varphi(g)\in\varphi(B^{G}_{m}(1))$ then $\ell(\varphi(g))\leq m$ by the triangle inequality. Since $\parallel g\parallel_{N}=\ell(\varphi(g))$ we have a contradiction.
$3\Rightarrow 4$. We preserve the notation of the previous paragraph. In particular $B^{G}_{m}(1)$ consists of all cosets from $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$. In order to prove that the set $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ is closed in $\mathsf{F}$ take any $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$. Thus $wN\not\in B^{G}_{m}(1)$. Choose a homomorphism $\varphi$ from $\mathsf{F}/N$ to some finite $H$ such that $\varphi(wN)\not\in\varphi(B^{G}_{m}(1))$. Then the corresponding homomorphism $\hat{\varphi}:\mathsf{F}\rightarrow\mathsf{F}/N\to H$ gives $\hat{\varphi}(w)\not\in\hat{\varphi}(B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N)$.
@@
Let $F_{1}$ be the set of all $g_{f,i}$ with $f\in F$ and $i\leq n_{f}$. We may assume that $F_{1}\subseteq F$ Then we associate $n_{f}=1$ to each $f\in F_{1}$.
Take a finite $Q\subset\mathbb{N}$. Let $n=\mathsf{max}(|f|\,|\,f\in F)$. We may assume that $Q$ is an initial segment of natural numbers and contains $\mathsf{max}\{\parallel fN\parallel_{N}\,|\,f\in F\cup B^{\mathsf{F},|\cdot|}_{2n}(1)\}+1$. Note that $n_{f}=\parallel f\parallel_{N}$ for each $f\in F$.
Since all $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$, $m\in Q$, are closed, for each $w\in(F\cup B^{\mathsf{F},|\cdot|}_{2n}(1))\setminus\{1\}$ there is a subgroup $H_{w}\triangleleft\,\mathsf{F}$ of finite index such that for every $m\in Q$ with $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ we have $wH_{w}\cap B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N=\emptyset$. Now it is easy to find a subgroup of finite index of $\mathsf{F}$, say $H_{F}$, such that for all $w\in(F\cup B^{\mathsf{F},|\cdot|}_{2n}(1))\setminus\{1\}$ and $m\in Q$ with $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ we have $wH_{F}\cap B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N=\emptyset$. In particular, taking $m=0$ we see $B^{\mathsf{F},|\cdot|}_{2n}(1)N\cap H_{F}\subseteq N$. The latter obviously means that the cosets from $B^{\mathsf{F},|\cdot|}_{n}(1)N\subset\mathsf{F}/N$ have pairwise distinct representatives in $\mathsf{F}/(H_{F}N)$.
Let us define a word norm on the finite group $\mathsf{F}/(H_{F}N)$ taking the cosets from $F^{\mathsf{F}}_{1}(H_{F}N)$ as a generating set. It can happen that $\langle F^{\mathsf{F}}_{1}(H_{F}N)\rangle\not=\mathsf{F}/(H_{F}N)$, i.e. the norm is defined only for elements of a proper subgroup of $\mathsf{F}/(H_{F}N)$. Then we just fix a sufficiently large natural number $\hat{m}$ (for example $\hat{m}>[\mathsf{F}:(H_{F}N)]$) and take it as the norm for the elements of $\mathsf{F}/(H_{F}N)\setminus\langle F^{\mathsf{F}}_{1}(H_{F}N)\rangle$. By the previous paragraph for each $g\in F\setminus\{1\}$ the norm of $g(H_{F}N)\in\mathsf{F}/(H_{F}N)$ defined in this way coincides with $n_{g}=\parallel g\parallel_{N}$. We see that the finite group $\mathsf{F}/(H_{F}N)$ has a norm witnessing residual finiteness of $\mathsf{F}/N$ as a metric group for the set $FN$.
@@
Note that under the homomorphism $\varphi$ the norm does not increase, i.e. it is a metric homomorphism. Indeed, if $\parallel h\parallel_{N}=m$, then $h\in B^{G}_{m}$, i.e. $\varphi(h)\in(\varphi(B^{G}_{1}))^{m}$. By the triangle inequality $\ell(\varphi(h))\leq m$. ∎

### v2

@@ 1 Introduction
Let $G$ be a group and let $\Lambda$ be a closed convex subset of $[0,\infty)$ so that $0\in\Lambda$ (the case $\Lambda=[0,\infty)$ is possible).
@@
If $\ell$ is a $\Lambda$-(pseudo)-norm on $G$ then we say that $(G,\ell)$ is a (pseudo) normed group. The pseudo-norm $\ell$ is invariant if $\ell(h^{-1}gh)=\ell(g)$ for all $g,h\in G$. In this case, it defines an invariant pseudo-metric by $d_{\ell}(g,h)=\ell(gh^{-1})$. Thus
@@
There are metric/normed groups for which the question of metric weak soficity looks very interesting. For example it would be interesting to know how it fits to bounded normal generation (with norms as in [7]), in particular in the case of Chevalley groups, see [18]. We also mention locally compact groups with property PL (i.e. unbounded length is proper), see [5]. Our original motivation is due to [6], [9], [14].
@@ 1.1 Word norms
Finitely generated groups with word norms are especially visible in (geometric) group theory, see [1], [3], [11], [12], [17] and [20]. For example when $G$ is generated by finitely many conjugacy classes then finiteness of its diameter for the appropriate word metric implies boundedness of its diameter with respect to any invariant metric, [1]. Thus metric weak soficity in the case of word norms is natural for investigations. This leads to an interesting subclass of abstract finitely generated weakly sofic groups: those ones which are metrically weakly sofic with respect to a word norm associated with some tuple of generators. It is an easy exercise that finitely generated abelian groups are metrically weakly sofic with respect to word norms arising from (Maltsev) bases. In fact they are metrically residually finite (See 2.3). There are other examples in [10] (which are metrically LEF). These results support the following problem:
@@
In order to approach this problem we give a characterization of the metric weak soficity in terms of profinite topology. The following statement is the main result of the paper.
Theorem 4.1. Let $\mathsf{F}$ be a finitely generated free group and $G=\mathsf{F}/N$. Assume that $S\subseteq\mathsf{F}$ is a finite symmetric set representing elements of $G$ that define an invariant word norm, say $\parallel\cdot\parallel_{N}$, in $G$.
The normed group $(G,\parallel\cdot\parallel_{N})$ is metrically weakly sofic if and only if for any finite sequence $g_{1}$, $g_{2}$,…, $g_{k}\in G$ of $\parallel\cdot\parallel_{N}$-norm 1 and $h_{1}$, $h_{2}$,…, $h_{l}$ from $N$ (possibly with repetitions), the closure of every product $g_{i_{1}}^{\mathsf{F}}\cdot g_{i_{2}}^{\mathsf{F}}\ldots\cdot g_{i_{m}}^{\mathsf{F}}\cdot h_{1}^{\mathsf{F}}\cdot h_{2}^{\mathsf{F}}\ldots h_{l}^{\mathsf{F}}$ is a subset of $\{w\in\mathsf{F}|\parallel wN\parallel_{N}\leq m\}N$.
@@
In Sections 3 and 5 we give theorems which characterize metric residual finiteness and metric LEF in terms of profinite topology. We also discuss a number of questions that arise in our approach.
@@ 2 Length functions on a group
This is a preliminary section. When we consider a normed group $(G,\ell)$, it is usually assumed that the norm is invariant. A free group with the standard length function ||$\cdot$|| ⋅ | with respect to a free basis will be the only exception to this rule. Just in case in Section 2.1 we do not make the general invariantness assumption.
@@ 2.1 Standard properties of norms
Then the family $\{B_{<r}(g)\,|\,r\in(\Lambda\cap\mathbb{Q})\setminus\{0\},\,g\in G\}$ is a subbase of the topology defined by $\ell$. Thus, every normed group is also a topological group.
@@ 2.3 Metric RF and metric LEF
It is worth noting that when in Definition 2.7 the values of $\ell_{G}$ or all possible $\ell_{C}$ belong to $\mathbb{Q}$ (in particular, when they are are integer-valued) the definition can be essentially simplified. The condition should be formulated as follows: for every finite $D\subset G$ there is a normed $(C,\ell_{C})\in\mathcal{C}$ and an injective map $\phi:D\rightarrow C$ such that
@@ 2.4 Norms for RF and metric LEF
It is easy to see that $\ell$ is an invariant norm which defines the profinite topology. In order to verify that $(G,\ell)$ is metrically fully residually finite take any finite $K\subset G$. Let $N_{s}$ be chosen so that $K\cap N_{s}\subseteq\{1\}$. It is easy to see that for every $g\in K$ we have $\ell(g)=\ell_{G/N_{s}}(gN_{s})$ where the right side norm is defined as in Lemma 2.2. ∎
@@
Let $\ell$ be an invariant norm on a group $G$ and $\mathsf{Rng}(\ell)\subseteq\Lambda_{0}$. Assume that $(G,\ell)$ is metrically LEF (resp. RF as a normed group).
Then $(G,\ell)$ is metrically LEF (resp. RF as a normed group) with respect to the subclass consisting of finite invariantly normed groups with norms having the set of values in $\Lambda_{0}$.
@@
Take a finite $K\subset G$, a finite $Q\subset\Lambda\cap\mathbb{Q}$ with $0\in Q$ and a $K$-$Q$-almost-homomorphism $\varphi$ from $(G,\ell)$ to some finite $(G_{0},\ell_{0})$ with an invariant norm. It can happen that $\mathsf{Rng}(\ell_{0})\not\subseteq\Lambda_{0}$. In order to correct this let us define a function $\ell^{\prime}$ on $G_{0}$ as follows. If $g\in G_{0}$ and $\ell_{0}(g)\in\Lambda_{0}$ put $\ell^{\prime}(g)=\ell_{0}(g)$. Note, that in this way we have $\ell(g)=\ell_{0}(g)$ for all $g\in\varphi(K)$, see Remark 2.8.
@@
In order to see property (iii) of Definition 1.1 let us apply induction. Let $m$ be the first number from $\Lambda_{0}$ such that $\ell_{0}$ takes values from $[0,m]\setminus\mathbb{N}$. Thus, these reals belong to the open interval $(m-1,m)$. Since all values of $\ell_{0}$ that do not accede to $m-1$, belong to $\mathbb{N}$, one easily sees that the (triangle) inequality (iii) holds for $\ell^{\prime}$ and for all $g$, $h$ with $\ell_{0}(gh)\leq m$.
@@
It seems to us that the method used in this proof can work in many other situations. In fact, we do not know any example in which the natural version of Proposition 2.10 does not hold. For example, it is easy to see that the statement of this proposition holds in the case when $\Lambda_{0}$ is dense in $\Lambda$.
3 Approximations of finitely generated normed groups with respect to word norms
@@ 3.1 Plan for Sections 3,4,5
Our aim is to describe the properties from the problem above in terms of the profinite topology of $\mathsf{F}$. In Section 3 we will study metric RF, in Section 4 - metric weak soficity and in Section 5 - metric LEF. In Section 4 we will show that $\mathsf{F}_{2}$ is not metrically weakly sofic as a metric group of the following form.
@@
In the remaining part of this introduction we give three useful remarks.
Remark 3.1
(see Section 2 of [8]) Closed sets in the profinite topology can be characterized by the following separability property: a set $X\subseteq G$ is closed in the profinite topology if and only if for any $g\not\in X$ there exists a homomorphism $\rho$ from $G$ to a finite group such that $\rho(g)\not\in\rho(X)$.
@@
This remark also holds for residual finiteness of normed groups where $\phi$ to be an abstract homomorphism from $G$ to a finite group.
In the third remark we again arrive at the the situation when $G=\langle\overline{S}\rangle$ with the corresponding word norm but now without assumption of finiteness of $S$. Note that the equalities $B_{1}(1)=\overline{S}\cup\{1\}$ and $B_{n}(1)=(B_{1}(1))^{n}$ always hold in such a group. We now notice that the converse also holds.
@@ 3.2 Metric RF
    The normed group $(G,\parallel\cdot\parallel_{N})$ is RF as a normed group;
@@
    The normed group $(G,\parallel\cdot\parallel_{N})$ is metrically RF;
@@
    For any $m\in\mathbb{N}$ and any $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ there is a homomorphism $\hat{\varphi}$ from $\mathsf{F}$ to a finite group such that $\hat{\varphi}(w)\not\in\hat{\varphi}(B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N)$ and $\hat{\varphi}(N)=1$.
Then $2\Rightarrow 3\Rightarrow 4\Rightarrow 1$, and assuming that $S$ is finite, $1\Leftrightarrow 2$.
@@
$2\Rightarrow 3$. W.l.o.g. we may assume that there are elements $g\in G$ with $\parallel g\parallel_{N}\geq 2$. In order to prove that the set $B^{G}_{m}(1)=\{x\in G\,|\,\parallel x\parallel_{N}\leq m\}$ is closed, take any $g\not\in B^{G}_{m}(1)$. Choose a metric homomorphism $\varphi$ from $G$ to some finite $(H,\ell)$ such that $\ell(\varphi(g))=\parallel g\parallel_{N}$. If $\varphi(g)\in\varphi(B^{G}_{m}(1))$ then $\ell(\varphi(g))\leq m$ by the triangle inequality. Since $\parallel g\parallel_{N}=\ell(\varphi(g))$ we have a contradiction.
$3\Rightarrow 4$. We preserve the notation of the previous paragraph. In particular $B^{G}_{m}(1)$ consists of all cosets from $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$. If $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$, then $wN\not\in B^{G}_{m}(1)$. Choose a homomorphism $\varphi$ from $\mathsf{F}/N$ to some finite $H$ such that $\varphi(wN)\not\in\varphi(B^{G}_{m}(1))$. Then the corresponding homomorphism $\hat{\varphi}:\mathsf{F}\rightarrow\mathsf{F}/N\to H$ gives $\hat{\varphi}(w)\not\in\hat{\varphi}(B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N)$ and $\hat{\varphi}(N)=1$.
@@
Let $F_{1}$ be the set of all $g_{f,i}$ with $f\in F$ and $i\leq n_{f}$. We may assume that $F_{1}\subseteq F$. Then we associate $n_{f}=1$ to each $f\in F_{1}$.
Take a finite $Q\subset\mathbb{N}$. Let $n=\mathsf{max}(|f|\,|\,f\in F)$. We may assume that $Q$ is an initial segment of natural numbers and contains $\mathsf{max}\{\parallel fN\parallel_{N}\,|\,f\in B^{\mathsf{F},|\cdot|}_{2n}(1)\}+1$. Note that $F\subseteq B^{\mathsf{F},|\cdot|}_{2n}(1)$, and $n_{f}=\parallel fN\parallel_{N}$ for each $f\in F$.
By the assumption of this part of the proof, for each $w\in B^{\mathsf{F},|\cdot|}_{2n}(1)\setminus\{1\}$ there is a subgroup $H_{w}\triangleleft\,\mathsf{F}$ of finite index such that $N\subseteq H_{w}$ and for every $m\in Q$ with $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ we have $wH_{w}\cap B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N=\emptyset$. Now we find a subgroup of finite index of $\mathsf{F}$, say $H_{F}$, such that $N\subseteq H_{F}$ and for all $w\in B^{\mathsf{F},|\cdot|}_{2n}(1)\setminus\{1\}$ and $m\in Q$ with $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ we have $wH_{F}\cap B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N=\emptyset$. In particular, taking $m=0$, we see $B^{\mathsf{F},|\cdot|}_{2n}(1)N\cap H_{F}\subseteq N$. The latter obviously means that the cosets from $B^{\mathsf{F},|\cdot|}_{n}(1)N\subset\mathsf{F}/N$ have pairwise distinct representatives in $\mathsf{F}/H_{F}$.
Let us define an invariant word norm on the finite group $\mathsf{F}/H_{F}$ taking the cosets from $F^{\mathsf{F}}_{1}H_{F}$ as a generating set. It can happen that $\langle F^{\mathsf{F}}_{1}H_{F}\rangle\not=\mathsf{F}/H_{F}$, i.e. the norm is defined only for elements of a proper subgroup of $\mathsf{F}/H_{F}$. Then we just fix a sufficiently large natural number $\hat{m}$ (for example $\hat{m}>[\mathsf{F}:H_{F}]$) and take it as the norm for the elements of $\mathsf{F}/H_{F}\setminus\langle F^{\mathsf{F}}_{1}H_{F}\rangle$. By the previous paragraph, for each $g\in F\setminus\{1\}$ the norm of $gH_{F}\in\mathsf{F}/H_{F}$ defined in this way coincides with $n_{g}=\parallel g\parallel_{N}$. We see that the finite group $\mathsf{F}/H_{F}$ has a norm that witnesses residual finiteness of $\mathsf{F}/N$ as a metric group for the set $FN$.
@@
Note that under the homomorphism $\varphi$ the norm does not increase, i.e. it is a metric homomorphism. Indeed, if $\parallel h\parallel_{N}=m$, then $h\in B^{G}_{m}(1)$, i.e. $\varphi(h)\in(\varphi(B^{G}_{1}(1)))^{m}$. By the triangle inequality $\ell(\varphi(h))\leq m$. ∎